### v1

@@ Abstract
Quantum computing has emerged as a transformative paradigm with the potential to outperform classical approaches on specialized computational tasks [1]. Concurrently, machine learning (ML) continues advancing rapidly, driven by increasing data availability and accelerated computing hardware [2]. Quantum machine learning (QML), at the intersection of these fields, has significant potential to unlock new capabilities in data processing and complex problem solving [2, 3]. By utilizing quantum phenomena such as superposition and entanglement, QML algorithms may effectively address high dimensionality and combinatorial complexity beyond classical counterparts [4, 5].
@@
Despite these prospects, large-scale quantum computing faces substantial practical challenges from noise and decoherence. Advanced error correction protocols can address these issues through sophisticated decoding algorithms, though their design complexity has led researchers to increasingly employ machine learning approaches for automation and enhancement of these decoding tasks [1]. Beyond error correction, QML applications have been explored extensively across domains such as image classification, natural language processing, and high energy physics [6, 7, 8, 9, 10]. Although proof of concept demonstrations on quantum processors like Google’s Sycamore [11] and IBM’s superconducting systems [12, 13] show feasibility, significant gaps persist between laboratory experiments and reliable industrial deployment [14, 15].
@@
A promising strategy for overcoming these gaps involves leveraging embedding and dimensionality reduction methods. Classical preprocessing techniques such as principal component analysis, or learned neural encoders like variational autoencoders, effectively reduce dataset complexity prior to quantum model input [8, 16]. These hybrid approaches help manage limited qubit resources on current quantum hardware while exploiting quantum enhanced feature spaces [17, 5]. Benchmark studies comparing quantum and classical ML approaches underline the potential and present limitations of QML [18, 14, 15]. Advancing this field requires systematic evaluations under realistic conditions, incorporating representative datasets, accurate noise models, and relevant performance metrics [2, 3, 19].
@@
We propose an embedding-aware, hybrid quantum-classical QSVM framework designed to address the scalability limitations of quantum machine learning. By integrating class-balanced -means data distillation with pretrained embeddings, our pipeline reduces data dimensionality while preserving task-relevant structure. Quantum kernel classification is performed using tensor network simulation with NVIDIA’s cuTensorNet [20, 21]. Benchmarking on MNIST and Fashion-MNIST shows that this embedding-driven approach consistently outperforms classical and quantum baselines in both accuracy and efficiency, confirming its value for scalable and resource-constrained quantum machine learning applications.
@@ 2.1 Scaling QML with Simulation Frameworks
Simulating larger circuits remains a popular strategy because near-term quantum devices have limited qubits. Efficient tensor-network methods can push simulations of quantum support vector machines (QSVMs) to hundreds of qubits [21], addressing scaling issues that plague naïve state-vector simulators. This line of research proves instrumental for prototyping advanced QML algorithms, guiding their eventual deployment on real hardware [16, 22].
@@ 2.2 Quantum Classifiers for Image Recognition
Variational quantum classifiers (VQCs), variational quantum circuits [6, 23, 22], quantum kernels, and hybrid architectures [7, 10, 3, 24, 25, 26] have been applied to standard benchmarks such as MNIST, Fashion-MNIST, and medical imaging tasks [6, 7, 9, 27]. Although classical deep networks often outperform small quantum models on large-scale datasets, quantum classifiers show competitive performance in data-scarce or high-dimensional settings by leveraging specialized embeddings and kernel methods [8, 26]. Several studies have also demonstrated end-to-end quantum classification pipelines executed on actual hardware, though typically limited to smaller datasets due to current device constraints [6, 10].
@@ 2.3 Hybrid Classical-Quantum Techniques
Several researchers propose hybrid approaches: classical layers for data preprocessing or encoding, followed by quantum layers for feature transformation or classification [7, 8, 16]. These methods can offset limited qubit counts by handing only compressed or task-relevant information to the quantum circuit [6]. Genetic algorithms [16], autoencoders [8], and transfer learning [10] have all been employed to optimize these hybrid models.
@@ 2.4 Applications Beyond Image Classification
QML has also been trialed in domains like high-energy physics [8, 13], medical diagnosis [7, 27], and scientific computing [21]. In certain specialized tasks - e.g., identifying the Higgs boson in proton collision data [8] - quantum models can match or exceed classical baselines under realistic noise models. These studies highlight the versatility of QML but also the pressing need for systematic benchmarking to compare cost-benefit trade-offs [18, 15].
As a whole, prior works illustrate QML’s broad applicability, from error decoding [1] and fundamental reviews [2, 14] to specialized classifiers for real-world tasks [6, 7, 10]. However, two critical gaps remain in the literature. First, a cohesive comparison that unifies insights across multiple domains and positions these results against robust classical baselines remains a key frontier [14, 15]. Second, existing approaches lack systematic investigation of how different embedding strategies affect quantum advantage, particularly the synergy between modern neural representations and quantum feature spaces.
@@ 3.1 Strategy Overview
Our approach addresses the scalability challenges of quantum machine learning through a hybrid quantum-classical pipeline that strategically combines data preprocessing, feature extraction, and quantum kernel methods. As illustrated in Figure 1, the framework operates through eight sequential stages: we begin with image data extraction and preprocessing, followed by data distillation using class-balanced -means clustering to reduce dataset size while maintaining representative samples. Next, we generate vector representations using pretrained models such as EfficientNet-B3 [28] and Vision Transformer variants [29], then apply Principal Component Analysis (PCA) to compress embeddings and match quantum hardware constraints. The processed embeddings are used to design a Quantum Support Vector Machine (QSVM) using the Tensor Network Support Matrix (TNSM) framework [21], which constructs quantum kernels through parameterized circuits and tensor network simulation using a data re-uploading and compute-uncompute strategy. Finally, we evaluate model performance through cross-validation and test on a held-out validation set to assess generalization capability. This embedding-aware strategy enables us to leverage the representational power of modern neural architectures while exploiting quantum kernel advantages for classification tasks, making quantum machine learning more practical and scalable for real-world applications.
@@ 3.2 Quantum Model Architecture
Chen et al. [21] forms the foundational simulation framework upon which our quantum model is built. The architecture employs a parameterized quantum circuit that encodes input data through rotational gates and entanglement layers within a Block-Encoded State (BPS) circuit, as shown in Figure 2. This data re-uploading circuit design has been validated in quantum learning applications and implemented within the Qiskit framework [30], serving as a proven foundation for kernel-based quantum classification.
@@
Implementation Optimizations. In contrast to the original implementation [21], we introduce several performance enhancements to the operand construction pipeline. First, to reduce redundant computations of trigonometric and exponential operations across repeated input angles, we apply function-level caching using Python’s @cache decorator. This memorization strategy significantly reduces overhead during the generation of gate matrices (e.g., parameterized and gates), which are frequently reused across multiple operand evaluations. Additionally, we precompute sine and cosine values in a shared utility function to avoid duplicating expressions and improve code modularity. Operand batches are generated via list comprehensions instead of iterative append calls, enhancing memory efficiency and readability. Likewise, tensor network amplitudes are computed using preallocated lists, eliminating dynamic resizing and reducing garbage collection pressure.
@@ 3.3 Quantum Kernel Formulation
The quantum kernel between two data points and is computed as the transition amplitude between their corresponding quantum states [3, 31]:
where represents the quantum feature map implemented by our parameterized circuit . The quantum advantage emerges from the exponentially large Hilbert space dimension compared to the classical feature space [4]. However, this advantage is critically dependent on how classical data is embedded before quantum encoding.
@@ 4.1 Dataset and Data Distillation Process
MNIST Dataset: We use the MNIST dataset [32], a widely recognized benchmark for evaluating image classification models. It consists of 70,000 grayscale images of handwritten digits (0–9), each of size pixels, divided into 60,000 training and 10,000 test samples.
Fashion-MNIST Dataset: We also utilize the Fashion-MNIST dataset [33], a benchmark dataset designed as a more challenging alternative to MNIST. It comprises 70,000 grayscale images of 10 fashion item categories (e.g., t-shirts, trousers, dresses), each of size pixels, split into 60,000 training and 10,000 test samples. This dataset provides a diverse set of visual patterns, enabling robust evaluation of our QSVM pipeline in a multi-class classification setting.
@@ 4.2 Embedding Extraction and Data Compression
To construct compact and informative inputs for quantum classification, we extract high-dimensional feature embeddings using pretrained image encoders. Specifically, we employ EfficientNet-B3 [28] and Vision Transformer (ViT) variants [29] trained via the CLIP framework [34]. These models, pretrained on large-scale datasets, capture rich semantic features that are well-suited for downstream classification tasks.
@@ 4.4 Computational Infrastructure
All experiments are conducted on NVIDIA A100 Tensor Core GPUs with 80GB HBM2 memory, using CUDA 12.0 and NVIDIA’s cuQuantum cuTensorNet backend [20] for quantum simulation. This high-performance computing environment ensures consistent benchmarking conditions and enables efficient tensor network contraction for quantum kernel computation. The GPU-accelerated simulation framework allows us to explore larger quantum circuits than would be feasible with CPU-based approaches.
@@ 7 Conclusion
We present an embedding-aware quantum-classical framework that systematically addresses scalability challenges in quantum machine learning by strategically combining class-balanced data distillation with pretrained embeddings. Building upon Chen et al.’s GPU-accelerated quantum kernel method [21], our pipeline successfully reduces computational complexity while achieving measurable performance improvements over classical baselines.
@@
Future work should pursue automated embedding and kernel selection strategies [35] to eliminate manual hyperparameter tuning, explore sophisticated dimensionality reduction techniques beyond PCA to better preserve semantic information, and develop optimized quantum circuit designs [36] for enhanced computational efficiency. Expanding empirical validation to medical imaging and other high-dimensional domains will prove critical for demonstrating broader practical utility.

### v2

@@ Abstract
Quantum computing has emerged as a transformative paradigm with the potential to outperform classical approaches on specialized computational tasks [bausch2024learning]. Concurrently, machine learning (ML) continues advancing rapidly, driven by increasing data availability and accelerated computing hardware [peral2024systematic]. Quantum machine learning (QML), at the intersection of these fields, has significant potential to unlock new capabilities in data processing and complex problem solving [peral2024systematic, havlivcek2019supervised]. By utilizing quantum phenomena such as superposition and entanglement, QML algorithms may effectively address high dimensionality and combinatorial complexity beyond classical counterparts [abbas2021power, senokosov2024quantum].
@@
Despite these prospects, large-scale quantum computing faces substantial practical challenges from noise and decoherence. Advanced error correction protocols can address these issues through sophisticated decoding algorithms, though their design complexity has led researchers to increasingly employ machine learning approaches for automation and enhancement of these decoding tasks [bausch2024learning]. Beyond error correction, QML applications have been explored extensively across domains such as image classification, natural language processing, and high energy physics [shen2024classification, don2024fusion, belis2024guided, vasques2023application, kim2023classical]. Although proof of concept demonstrations on quantum processors like Google’s Sycamore [huang2022quantum] and IBM’s superconducting systems [gentinetta2024complexity, cugini2023comparing] show feasibility, significant gaps persist between laboratory experiments and reliable industrial deployment [gujju2024quantum, basilewitsch2024quantum].
@@
A promising strategy for overcoming these gaps involves leveraging embedding and dimensionality reduction methods. Classical preprocessing techniques such as principal component analysis, or learned neural encoders like variational autoencoders, effectively reduce dataset complexity prior to quantum model input [belis2024guided, phalak2024optimizing]. These hybrid approaches help manage limited qubit resources on current quantum hardware while exploiting quantum enhanced feature spaces [beer2020training, senokosov2024quantum]. Benchmark studies comparing quantum and classical ML approaches underline the potential and present limitations of QML [havenstein2018comparisons, gujju2024quantum, basilewitsch2024quantum]. Advancing this field requires systematic evaluations under realistic conditions, incorporating representative datasets, accurate noise models, and relevant performance metrics [peral2024systematic, havlivcek2019supervised, potempa2022comparing].
@@
We propose an embedding-aware, hybrid quantum-classical QSVM framework designed to address the scalability limitations of quantum machine learning. By integrating class-balanced -means data distillation with pretrained embeddings, our pipeline reduces data dimensionality while preserving task-relevant structure. Quantum kernel classification is performed using tensor network simulation with NVIDIA’s cuTensorNet [cutensornet, chen2024validating]. Benchmarking on MNIST and Fashion-MNIST shows that this embedding-driven approach consistently outperforms classical and quantum baselines in both accuracy and efficiency, confirming its value for scalable and resource-constrained quantum machine learning applications.
@@ 2.1 Scaling QML with Simulation Frameworks
Simulating larger circuits remains a popular strategy because near-term quantum devices have limited qubits. Efficient tensor-network methods can push simulations of quantum support vector machines (QSVMs) to hundreds of qubits [chen2024validating], addressing scaling issues that plague naïve state-vector simulators. This line of research proves instrumental for prototyping advanced QML algorithms, guiding their eventual deployment on real hardware [phalak2024optimizing, mitarai2018quantum].
@@ 2.2 Quantum Classifiers for Image Recognition
Variational quantum classifiers (VQCs), variational quantum circuits [shen2024classification, farhi2018classification, mitarai2018quantum], quantum kernels, and hybrid architectures [don2024fusion, kim2023classical, havlivcek2019supervised, sharma2023role, chen2022generating, blank2020quantum] have been applied to standard benchmarks such as MNIST, Fashion-MNIST, and medical imaging tasks [shen2024classification, don2024fusion, vasques2023application, maouaki2024quantum]. Although classical deep networks often outperform small quantum models on large-scale datasets, quantum classifiers show competitive performance in data-scarce or high-dimensional settings by leveraging specialized embeddings and kernel methods [belis2024guided, blank2020quantum]. Several studies have also demonstrated end-to-end quantum classification pipelines executed on actual hardware, though typically limited to smaller datasets due to current device constraints [shen2024classification, kim2023classical].
@@ 2.3 Hybrid Classical-Quantum Techniques
Several researchers propose hybrid approaches: classical layers for data preprocessing or encoding, followed by quantum layers for feature transformation or classification [don2024fusion, belis2024guided, phalak2024optimizing]. These methods can offset limited qubit counts by handing only compressed or task-relevant information to the quantum circuit [shen2024classification]. Genetic algorithms [phalak2024optimizing], autoencoders [belis2024guided], and transfer learning [kim2023classical] have all been employed to optimize these hybrid models.
@@ 2.4 Applications Beyond Image Classification
QML has also been trialed in domains like high-energy physics [belis2024guided, cugini2023comparing], medical diagnosis [don2024fusion, maouaki2024quantum], and scientific computing [chen2024validating]. In certain specialized tasks - e.g., identifying the Higgs boson in proton collision data [belis2024guided] - quantum models can match or exceed classical baselines under realistic noise models. These studies highlight the versatility of QML but also the pressing need for systematic benchmarking to compare cost-benefit trade-offs [havenstein2018comparisons, basilewitsch2024quantum].
As a whole, prior works illustrate QML’s broad applicability, from error decoding [bausch2024learning] and fundamental reviews [peral2024systematic, gujju2024quantum] to specialized classifiers for real-world tasks [shen2024classification, don2024fusion, kim2023classical]. However, two critical gaps remain in the literature. First, a cohesive comparison that unifies insights across multiple domains and positions these results against robust classical baselines remains a key frontier [gujju2024quantum, basilewitsch2024quantum]. Second, existing approaches lack systematic investigation of how different embedding strategies affect quantum advantage, particularly the synergy between modern neural representations and quantum feature spaces.
@@ 3.1 Strategy Overview
Our approach addresses the scalability challenges of quantum machine learning through a hybrid quantum-classical pipeline that strategically combines data preprocessing, feature extraction, and quantum kernel methods. As illustrated in Figure 1, the framework operates through eight sequential stages: we begin with image data extraction and preprocessing, followed by data distillation using class-balanced -means clustering to reduce dataset size while maintaining representative samples. Next, we generate vector representations using pretrained models such as EfficientNet-B3 [tan2019efficientnet] and Vision Transformer variants [dosovitskiy2020image], then apply Principal Component Analysis (PCA) to compress embeddings and match quantum hardware constraints. The processed embeddings are used to design a Quantum Support Vector Machine (QSVM) using the Tensor Network Support Matrix (TNSM) framework [chen2024validating], which constructs quantum kernels through parameterized circuits and tensor network simulation using a data re-uploading and compute-uncompute strategy. Finally, we evaluate model performance through cross-validation and test on a held-out validation set to assess generalization capability. This embedding-aware strategy enables us to leverage the representational power of modern neural architectures while exploiting quantum kernel advantages for classification tasks, making quantum machine learning more practical and scalable for real-world applications.
@@ 3.2 Quantum Model Architecture
Chen et al. [chen2024validating] forms the foundational simulation framework upon which our quantum model is built. The architecture employs a parameterized quantum circuit that encodes input data through rotational gates and entanglement layers within a Block-Encoded State (BPS) circuit, as shown in Figure 2. This data re-uploading circuit design has been validated in quantum learning applications and implemented within the Qiskit framework [javadi2024quantum], serving as a proven foundation for kernel-based quantum classification.
@@
Implementation Optimizations. In contrast to the original implementation [chen2024validating], we introduce several performance enhancements to the operand construction pipeline. First, to reduce redundant computations of trigonometric and exponential operations across repeated input angles, we apply function-level caching using Python’s @cache decorator. This memorization strategy significantly reduces overhead during the generation of gate matrices (e.g., parameterized and gates), which are frequently reused across multiple operand evaluations. Additionally, we precompute sine and cosine values in a shared utility function to avoid duplicating expressions and improve code modularity. Operand batches are generated via list comprehensions instead of iterative append calls, enhancing memory efficiency and readability. Likewise, tensor network amplitudes are computed using preallocated lists, eliminating dynamic resizing and reducing garbage collection pressure.
@@ 3.3 Quantum Kernel Formulation
The quantum kernel between two data points and is computed as the transition amplitude between their corresponding quantum states [havlivcek2019supervised, schuld2019quantum]:
where represents the quantum feature map implemented by our parameterized circuit . The quantum advantage emerges from the exponentially large Hilbert space dimension compared to the classical feature space [abbas2021power]. However, this advantage is critically dependent on how classical data is embedded before quantum encoding.
@@ 4.1 Dataset and Data Distillation Process
MNIST Dataset: We use the MNIST dataset [lecun1998gradient], a widely recognized benchmark for evaluating image classification models. It consists of 70,000 grayscale images of handwritten digits (0–9), each of size pixels, divided into 60,000 training and 10,000 test samples.
Fashion-MNIST Dataset: We also utilize the Fashion-MNIST dataset [xiao2017fashion], a benchmark dataset designed as a more challenging alternative to MNIST. It comprises 70,000 grayscale images of 10 fashion item categories (e.g., t-shirts, trousers, dresses), each of size pixels, split into 60,000 training and 10,000 test samples. This dataset provides a diverse set of visual patterns, enabling robust evaluation of our QSVM pipeline in a multi-class classification setting.
@@ 4.2 Embedding Extraction and Data Compression
To construct compact and informative inputs for quantum classification, we extract high-dimensional feature embeddings using pretrained image encoders. Specifically, we employ EfficientNet-B3 [tan2019efficientnet] and Vision Transformer (ViT) variants [dosovitskiy2020image] trained via the CLIP framework [radford2021learning]. These models, pretrained on large-scale datasets, capture rich semantic features that are well-suited for downstream classification tasks.
@@ 4.4 Computational Infrastructure
All experiments are conducted on NVIDIA A100 Tensor Core GPUs with 80GB HBM2 memory, using CUDA 12.0 and NVIDIA’s cuQuantum cuTensorNet backend [cutensornet] for quantum simulation. This high-performance computing environment ensures consistent benchmarking conditions and enables efficient tensor network contraction for quantum kernel computation. The GPU-accelerated simulation framework allows us to explore larger quantum circuits than would be feasible with CPU-based approaches.
@@ 7 Conclusion
We present an embedding-aware quantum-classical framework that systematically addresses scalability challenges in quantum machine learning by strategically combining class-balanced data distillation with pretrained embeddings. Building upon Chen et al.’s GPU-accelerated quantum kernel method [chen2024validating], our pipeline successfully reduces computational complexity while achieving measurable performance improvements over classical baselines.
@@
Future work should pursue automated embedding and kernel selection strategies [Incudini_2024] to eliminate manual hyperparameter tuning, explore sophisticated dimensionality reduction techniques beyond PCA to better preserve semantic information, and develop optimized quantum circuit designs [sünkel2023ga4qcogeneticalgorithmquantum] for enhanced computational efficiency. Expanding empirical validation to medical imaging and other high-dimensional domains will prove critical for demonstrating broader practical utility.